18

18-7Codeine (C_18H_21NO_3) is a narcotic pain reliever
Answer #1

Pkb = 5.80

-logKb = 5.80

Kb = 1.585*10^-6

Ka*Kb = 10^-14

Ka = 10^-14/(1.585*10^-6) = 6.31*10^-9

[H+] =(Ka*C)^1/2

[H+] = (6.31*10^-9*0.036)^1/2 = 1.507*10^-10

Ph = -logH+] = -log[1.507*10^-10]

Ph = 9.822